OpenStudy (anonymous):

Carlon plans to evaporate a number of gallons of water from 90 gallons of a 15% salt solution to make a 25% salt solution. Which equation can he use to find x, the number of gallons that should be evaporated?

OpenStudy (australopithecus):

Do you know how to set up ratios?

OpenStudy (anonymous):

i don't think so..

OpenStudy (anonymous):

the options are 1. 13.5/90-x = 25/100 2. 13.5/x =25/100 3. 13.5(90-x) = 0.25 4. 13.5(90-x) = 25

OpenStudy (australopithecus):

Use the formula Volume Solute/(Volume Solute + volume of solvent) = percent/100

OpenStudy (australopithecus):

Remember you are only removing water from the solution not whatever isn't water at least this is what I assume

OpenStudy (australopithecus):

I guess we shouldnt say the compound is a volume we should just give it arbitrary untis

OpenStudy (australopithecus):

You can determine how much 15% of 90 gallons makes up. With that new number and the percentage you are aiming for you can figure out the gallons of water needed to be removed

OpenStudy (australopithecus):

You can then just subtract the aimed for volume from the old volume to figure out how much you want to evaporate. Do you follow? Damn I'm tired why did this take me so long to figure out blah I need to eat or something probably who knows I know I dont

OpenStudy (anonymous):

yeah i get it, and sorry..

OpenStudy (australopithecus):

dont be sorry :D

OpenStudy (australopithecus):

if you tell me your answer I can confirm it for you

OpenStudy (anonymous):

i got 13.5(90-x)=0.25??

OpenStudy (australopithecus):

x = 89.9815ga for that, that isnt a reasonable answer

OpenStudy (anonymous):

ohh.. ok

OpenStudy (australopithecus):

do you understand percentages?

OpenStudy (anonymous):

yeah..

OpenStudy (australopithecus):

I'm not trying to be condescending, so please dont take offense

OpenStudy (anonymous):

ok.

OpenStudy (australopithecus):

-the way you should go about these problems is first convert over from percentages to decimals, percentages are just used because they are pretty they dont have much function other than that (people dont like looking at decimals) oops sorry I'm out of it right now ha, i see your reasoning wasnt based on your lack of understanding precentages I just read what your answer was incorrectly (it is still incorrect none the less) -Next as I pointed out the water is changing in amount but the salt is not, so we should find the amount of salt present in the solution 0.15(90) = 13.5 Using 13.5 we can figure out the final volume we want 13.5/y = 0.25 where y is the desired final volume. so, 90 - y = x where x is the amount we want out of solution Use system of equations to solve Using system of equations seems pretty unnecessary for this problem but they probably wanted to see if you could express this problem in one equation

OpenStudy (australopithecus):

sorry for taking so long to answer the question, I need to sleep.

OpenStudy (australopithecus):

and possibly offending you I thought you forgot to convert a percentage in your choice of answer because I'm durpy right now

OpenStudy (australopithecus):

I guess you figured out the answer yourself? Anyways if you have any questions about my solution please feel free to ask. If you need a definition for system of equations I also do not mind explaining it

OpenStudy (anonymous):

no problem, and thank you